# Quantum Computational Insurance and Actuarial Science

### v1

@@ VI Conclusion and discussion
In the preceding sections, we delved into the realm of quantum computational insurance and actuarial science. We examined critical issues across the principal branches of insurance, including non-life insurance, life insurance, and reinsurance. Following this, we explored the application of quantum algorithms to address these challenges. Additionally, we conducted both numerical and experimental studies on several insurance problems using quantum algorithms. In this section, we outline the roadmap for the future of quantum computational insurance.
@@
In the NISQ era, the available qubit number of quantum operation depth is limited due to the noise [20]. Although experiments have shown early quantum advantages on some specific applications [65], there is still a long way to universal quantum computing. While one of the most challenging problems is to strive for fault-tolerant quantum computing, how can the insurance industry make use of quantum computation in the NISQ era? We believe that both theoretical and experimental breakthroughs focusing on quantum insurance and actuarial science will be achieved. On the aspect of theoretical development, many quantum insurance algorithms will be proposed and proved to have quantum speed-up. Quantum stochastic modeling will be applied to study risk and ruin theory with many applications as suggested in the previous section. Also, quantum machine learning and optimization will be further developed on prediction and portfolio optimization problems. As for the aspect of experiments, quantum insurance algorithms, including but not limited to variational quantum algorithms and quantum Monte Carlo simulations, can be performed on quantum processors. The insurance industry can help to find the best scenarios and applications of quantum insurance and bring more attention to this cross field. With the cooperation of quantum computation scientists the insurance industry and academia, quantum algorithms driven or inspired by insurance can reveal more power of quantum computation and further applicability to insurance.
@@
VII Methods
VII.1 Quantum excess evaluation algorithm
Here, we provide details about the quantum excess evaluation algorithm. The algorithm is divided into the following steps:
State preparation: When $n$ qubits are used to prepare the distribution of loss, we discrete the domain $[0,10]$ with $N=2^{n}$ points. Then the state is prepared as:
$|\psi_{1}\rangle=\sum_{j=0}^{N-1}\sqrt{p_{j}}|j\rangle,$ (17)
where $p_{j}=f(X_{j})/C$ with $C$ the normalization factor.
Quantum subtractor According to the definition of the payment function, we first need to evaluate $x-M(x)=0.4(x-1)$ when $x>1$. which needs to evaluate $x-1$ with a quantum Subtractor detailed in [77]. Since in the $n$-qubit case, the loss range 10 is divided into $12^{n}-$ 1 parts, then the threshold 1 can be represented by a n-bit binary representation as $1_{n}=\text{bin}(\frac{2^{n}-1}{10})$. With one auxiliary qubit, we can finish this task and the resulting state is:
$|\psi_{2}\rangle=\sum_{j=0}^{1_{n}-1}\sqrt{p_{j}}|N-j-1_{n}\rangle|1\rangle+\sum_{j=1_{n}}^{N-1}\sqrt{p_{j}}|j-1_{n}\rangle|0\rangle.$ (18)
Controlled rotations Apply a $X$ gate on the last qubit, and then add one new measurement qubit, followed by a $RY(\pi/2)$ gate, we have:
$\sum_{j=0}^{1_{n}-1}\sqrt{p_{j}}|N-x-1_{n}\rangle|0\rangle\left[\cos(\frac{\pi}{4})|0\rangle+\sin(\frac{\pi}{4})|1\rangle\right]$ (19)
$+$ $\sum_{j=1_{n}}^{N-1}|j-1_{n}\rangle|1\rangle\left[\cos(\frac{\pi}{4})|0\rangle+\sin(\frac{\pi}{4})|1\rangle\right].$
Then apply sequences of $CCRY$ gate, we get:
$\sum_{x=0}^{1_{n}-1}{\sqrt{p_{x}}|N-x-1_{n}\rangle|0\rangle}{(\cos\frac{\pi}{4}|0\rangle+\sin\frac{\pi}{4}|1\rangle)}$ (20)
$+$ $\sum_{x=1_{n}}^{N-1}{\sqrt{p_{x}}|x-1_{n}\rangle|1\rangle}\otimes(\cos\theta|0\rangle+\sin\theta|1\rangle),$
where $\theta=\frac{\pi}{4}+0.4c(x-1_{n})$ and $c$ is a small number.
Measure and post-processing Finally, measure the measurement qubit and the probability of obtaining $|1\rangle$ is
$\sum_{x=0}^{1_{n}-1}{p_{x}}{\sin^{2}\frac{\pi}{4}}+\sum_{x=1_{n}}^{N-1}{p_{x}}{\sin^{2}(\frac{\pi}{4}+0.4c(x-1_{n}))}$ (21)
$\approx$ $\frac{1}{2}+\sum_{x=1_{n}}^{N-1}{0.4c{p_{x}}(x-1_{n})}$
After some simple processing, we can get the desired result. The quantum circuit is shown in Fig. 5.
@@ VII.1 Quantum excess evaluation algorithm
[…]
$\bm{\vdots}$
[…]
@@
Figure 5: Quantum circuits for the quantum excess evaluation algorithm. The unitary operations $U_{X}$ and $U_{S}$ are used to prepare the state $|\psi_{1}\rangle$ and realize the subtraction operations. Followed controlling operations rotate the measurement qubit. The desired value is associated with the probability of obtaining $|1\rangle$.
VII.2 Quantum singular value decomposition algorithm
@@
In this part, we briefly introduce the QSVD algorithm, which is detailed in [63]. The SVD of a matrix $D$ is $D=\sum_{i}\sigma_{i}|u_{i}\rangle\langle v_{i}|$, where $\sigma_{i}$ is the singular value, $|u_{i}\rangle/|v_{i}\rangle$ is the associated left/right singular vector. The QSVD algorithm starts with a vectorized representation of the matrix, which uses a vector of $MN$ elements to represent the $M\times N$ matrix.
$D$ $=\begin{pmatrix}D_{11}&\cdots&D_{1N}\\ \vdots&&\vdots\\ D_{M1}&\cdots&D_{MM}\end{pmatrix}$ (22)
$\to|D\rangle$ $=\begin{pmatrix}D_{11}&\cdots&D_{1N}&\cdots&D_{M1}&\cdots&D_{MM}\end{pmatrix}^{\operatorname{T}}.$
Therefore, the SVD of a matrix is transformed as
@@
In quantum computers, with a re-normalization factor $C=\sqrt{\langle D|D\rangle}$ we can prepare the quantum state of $D$ in the Schmidt decomposition form
@@
where $A,B$ represent the two-subsystems. Consider two unitary operations $U$ and $V$ that respectively convert the state $\{|u_{i}\rangle\}$ and $\{|v_{i}\rangle\}$ to the computational basis $\{|e_{i}\rangle\}$:
@@
Therefore, when measuring the two subsystems $A$ and $B$ of the state $|\psi\rangle_{AB}$, we will get the same measurement result in the computational basis. The corresponding probability is $\lambda_{i}^{2}=\sigma_{i}^{2}/C^{2}$. Finally, apply $U^{\dagger}$ and $V^{\dagger}$ on the computational basis state $|e_{i}\rangle$ will result in the singular vector $|u_{i}\rangle$ and $|v_{i}\rangle$.
To obtain $U/V$. we apply the variational QSVD algorithm by introducing an optimization procedure. That is: after preparing the state $|\psi\rangle_{AB}$, we apply two parameterized quantum circuits $U_{A}(\bm{\theta}_{1})$ and $V_{B}(\bm{\theta}_{2})$. The loss function is
@@
where $N_{m}$ is the number of shots, $p_{x},q_{x}$ are the two binary strings at the $x$-th shot. $\delta(a,b)$ is 0 if $a\neq b$ and 1 if $a=b$. A classical optimizer is employed to optimize the parameters to minimize the loss function.
ACKNOWLEDGEMENT
This work has been supported by the National Key Research and Development Program of China (Grant No. 2023YFB4502500), the National Natural Science Foundation of China (Grant No. 12404564), and the Aeronautical Science Foundation of China (Grant No. 2022Z073004001).

### v2

@@ 5 Conclusion and discussion
In the preceding sections, we delved into the realm of quantum computational insurance and actuarial science. We examined critical issues across the principal branches of insurance, including non-life insurance, life insurance, and reinsurance. Following this, we explored the application of quantum algorithms to address these challenges. Additionally, we conducted both numerical and experimental studies on several insurance problems using quantum algorithms. Besides the introduction of the general paths to applying quantum computing to insurance, these demonstrations clarified the associated technical details. The connection between quantum computation and the insurance industry is established and enhanced, which can promote the development of both quantum computing and insurance.
@@
In this section, we outline the roadmap for the future of quantum computational insurance. In the NISQ era, the available qubit number of quantum operation depth is limited due to the noise [42]. Although experiments have shown early quantum advantages on some specific applications [85], there is still a long way to universal quantum computing. While one of the most challenging problems is to strive for fault-tolerant quantum computing, how can the insurance industry make use of quantum computation in the NISQ era? We believe that both theoretical and experimental breakthroughs focusing on quantum insurance and actuarial science will be achieved. On the aspect of theoretical development, many quantum insurance algorithms will be proposed and proved to have quantum speed-up. Quantum stochastic modeling will be applied to study risk and ruin theory with many applications as suggested in the previous section. Also, quantum machine learning and optimization will be further developed on prediction and portfolio optimization problems. As for the aspect of experiments, quantum insurance algorithms, including but not limited to variational quantum algorithms and quantum Monte Carlo simulations, can be performed on quantum processors. The insurance industry can help to find the best scenarios and applications of quantum insurance and bring more attention to this cross-field. With the cooperation of quantum computation scientists, the insurance industry, and academia, quantum algorithms driven or inspired by insurance can reveal more power of quantum computation and further applicability to insurance.
@@
Acknowledgements This work has been supported by the National Key Research and Development Program of China (Grant No. 2023YFB4502500), the National Natural Science Foundation of China (Grant No. 12404564), and the Aeronautical Science Foundation of China (Grant No. 2022Z073004001).
@@
Data Availability The data that support the findings of this study are available from the corresponding author upon reasonable request.
Appendix A Quantum excess evaluation algorithm
Here, we provide details about the quantum excess evaluation algorithm. The algorithm is divided into the following steps, and a sketch map of the quantum circuit is shown in Fig. 6.
Setup: As introduced in the main text, the algorithm aims to estimate $\sum_{X_{j}\geq 1}f(X_{j})(X_{j}-M(X_{j}))$. We start this with representing $x$ with $n$ qubits $|x\rangle=|x_{n-1}\cdots x_{1}x_{0}\rangle$ under the condition that $x=\sum_{i=0}^{n-1}x_{i}2^{i}$, where $x_{i}\in\{0,1\},\forall i$. Then we re-scale the interval $[0,10]$ to $0,2^{n}-1]$ such that the state $|x\rangle$ indeed represents the number $10x/(2^{n}-1)$ (For instance, $|0\rangle^{\otimes n}$ and $|1\rangle^{\otimes n}$ refer to 0 and 10, respectively).
State preparation: We use $n$ qubits to represent each loss $X_{j}$ as $|j\rangle$. Then, using quantum superposition, we prepare the quantum state
$|\psi_{1}\rangle=\sum_{j=0}^{N-1}\sqrt{p_{j}}\,|j\rangle$ (19)
with the unitary $U_{X}$, where $p_{j}=f(X_{j})/C$ with $C$ a normalization factor.
Quantum subtractor: When $x\geq 1$, the payment is xM(x)=0.4(x1)x-M(x)=0.4(x-1). We first need to evaluate x1x-1 with a quantum Subtractor ($U_{S}$) detailed in [97], in which the threshold 1 needs to be converted as $1_{n}=\text{bin}(\tfrac{2^{n}-1}{10})$. With one auxiliary qubit, we can finish this task and the resulting state is:
$|\psi_{2}\rangle=\sum_{j=0}^{1_{n}-1}\sqrt{p_{j}}\,|N-j-1_{n}\rangle|1\rangle_{a}+\sum_{j=1_{n}}^{N-1}\sqrt{p_{j}}\,|j-1_{n}\rangle|0\rangle_{a},$ (20)
where $q_{a}$ means the auxiliary qubit.
Controlled rotations: Apply an $X$ gate on the auxiliary qubit, and then add one new measurement qubit $q_{m}$, followed by an $RY(\pi/2)$ gate; we have:
$|\psi_{3}\rangle=$ $\sum_{j=0}^{1_{n}-1}\sqrt{p_{j}}\,|N-x-1_{n}\rangle|0\rangle_{a}\left[\cos\!\bigl(\tfrac{\pi}{4}\bigr)|0\rangle_{m}+\sin\!\bigl(\tfrac{\pi}{4}\bigr)|1\rangle_{m}\right]$ (21)
$+\sum_{j=1_{n}}^{N-1}\sqrt{p_{j}}\,|j-1_{n}\rangle|1\rangle_{a}\left[\cos\!\bigl(\tfrac{\pi}{4}\bigr)|0\rangle_{m}+\sin\!\bigl(\tfrac{\pi}{4}\bigr)|1\rangle_{m}\right].$
Consider the $CCRY(\theta)$ gate, which has two control qubits and one target qubit. It applies an $RY(\theta)$ gate on the target qubit only if the two control qubits are both in the state $|1\rangle$:
$CCRY(\theta)\,|1\rangle|1\rangle|\psi\rangle=|1\rangle|1\rangle\,RY(\theta)\,|\psi\rangle,\qquad CCRY(\theta)\,|1\rangle|0\rangle|\psi\rangle=|1\rangle|0\rangle|\psi\rangle.$ (22)
Then using $RY(0)=I$, we can combine the above two equations as
$CCRY(\theta)\,|1\rangle|x\rangle|\psi\rangle=|1\rangle|x\rangle\,RY(\theta x)\,|\psi\rangle.$ (23)
Then we apply a sequence of $n$ CCRYCCRY gates shown in Fig. 6. For each gate, the first control qubit is the auxiliary qubit $q_{a}$. The second control qubit is one of the $n$ qubits that prepare $|j\rangle$. The rotation angle is $2^{i}c$, where $c$ is a relatively small positive number and $i$ is the index in the binary representation. According to Eq. (23), these operations will convert the state $|x_{n-1}\rangle\cdots|x_{1}\rangle|x_{0}\rangle|1\rangle|\psi\rangle$ to $|x_{n-1}\rangle\cdots|x_{1}\rangle|x_{0}\rangle|1\rangle RY\!\bigl(c\!\sum_{i=0}^{n-1}2^{i}x_{i}\bigr)\,|\psi\rangle$, or more accurately, $|x_{n-1}\rangle\cdots|x_{1}\rangle|x_{0}\rangle|1\rangle RY(cx)\,|\psi\rangle$. Therefore, $|\psi_{3}\rangle$ is evolved to the following state:
$|\psi_{4}\rangle=$ $\sum_{x=0}^{1_{n}-1}\sqrt{p_{x}}\;|N-x-1_{n}\rangle|0\rangle_{a}\bigl(\cos\tfrac{\pi}{4}|0\rangle_{m}+\sin\tfrac{\pi}{4}|1\rangle_{m}\bigr)$ (24)
$+\sum_{x=1_{n}}^{N-1}\sqrt{p_{x}}\;|x-1_{n}\rangle|1\rangle_{a}\otimes\bigl(\cos\theta_{x}|0\rangle_{m}+\sin\theta_{x}|1\rangle_{m}\bigr),$
where $\theta_{x}=\tfrac{\pi}{4}+c(x-1_{n})$.
Measure and post-processing: Finally, we measure the measurement qubit $q_{m}$, and the probability of obtaining $|1\rangle$ is
$P_{0}=\sum_{x=0}^{1_{n}-1}p_{x}\,\sin^{2}\!\bigl(\tfrac{\pi}{4}\bigr)+\sum_{x=1_{n}}^{N-1}p_{x}\,\sin^{2}\!\bigl(\tfrac{\pi}{4}+c(x-1_{n})\bigr).$ (25)
When $c$ is very small, we use Taylor’s expansion and neglect terms of order $\mathcal{O}(c^{2})$; the result is $\sin^{2}\!\bigl(\tfrac{\pi}{4}+c(x-1_{n})\bigr)\approx c(x-1_{n})+\tfrac{1}{2}$. Finally, the above probability is:
$P_{0}=\sum_{x=0}^{1_{n}-1}\tfrac{1}{2}p_{x}+\sum_{x=1_{n}}^{N-1}p_{x}\Bigl(\tfrac{1}{2}+c(x-1_{n})\Bigr)=\tfrac{1}{2}\sum_{x=0}^{N-1}p_{x}+c\sum_{x=1_{n}}^{N-1}p_{x}(x-1_{n}),$ (26)
With $\sum_{x}p_{x}=1$, it follows that:
$\sum_{x=1_{n}}^{N-1}p_{x}(x-1_{n})=\frac{P_{0}-\tfrac{1}{2}}{c}.$ (27)
We can see that the left-hand side is exactly what we need:
$\sum_{X_{j}\geq 1}f(X_{j})\bigl(X_{j}-M(X_{j})\bigr)=0.4\times\frac{10}{2^{n}-1}\sum_{x=1_{n}}^{N-1}p_{x}(x-1_{n}).$ (28)
Then the algorithm is finished.
@@ Appendix A Quantum excess evaluation algorithm
Figure 6: Quantum circuits for the quantum excess evaluation algorithm. The unitary operations $U_{X}$ and $U_{S}$ are used to prepare the state $|\psi_{1}\rangle$ and realize the subtraction operations. Followed controlling operations rotate the measurement qubit. The desired value is associated with the probability of obtaining $|1\rangle$ in multiple shots.
Appendix B Quantum singular value decomposition algorithm
In this part, we briefly introduce the QSVD algorithm, which is detailed in [72].
Vectorized SVD: The SVD of a matrix $D$ is
$D=\sum_{i}\sigma_{i}|u_{i}\rangle\langle v_{i}|,$ (29)
@@ Appendix B Quantum singular value decomposition algorithm
where $\sigma_{i}$ is the singular value, and $|u_{i}\rangle$ or $|v_{i}\rangle$ is the associated left or right singular vector. We can also express it using a vectorized matrix representation, which employs a vector of $MN$ elements to represent the $M\times N$ matrix:
$D=\begin{pmatrix}D_{11}&\cdots&D_{1N}\\ \vdots&&\vdots\\ D_{M1}&\cdots&D_{MM}\end{pmatrix}\;\to\;|D\rangle=\begin{pmatrix}D_{11}&\cdots&D_{1N}&\cdots&D_{M1}&\cdots&D_{MM}\end{pmatrix}^{\mathrm{T}}.$ (30)
Therefore, the SVD of a vectorized matrix is expressed as
@@
QSVD algorithm: In quantum computers, with a re-normalization factor $C=\sqrt{\langle D|D\rangle}$, we can prepare the quantum state of $D$ and express it in the Schmidt decomposition form
@@
where $A,B$ represent the two subsystems. Consider two unitary operations $U$ and $V$ that convert the states $\{|u_{i}\rangle\}$ and $\{|v_{i}\rangle\}$ to the computational basis $\{|e_{i}\rangle\}$:
@@
Therefore, when measuring the two subsystems $A$ and $B$ of the state $|\psi\rangle_{AB}$, we obtain the same measurement result on the computational basis. The corresponding probability is $\lambda_{i}^{2}=\sigma_{i}^{2}/C^{2}$. Finally, applying $U^{\dagger}$ and $V^{\dagger}$ on the computational basis state $|e_{i}\rangle$ yields the singular vectors $|u_{i}\rangle$ and $|v_{i}\rangle$.
To obtain $U$ and $V$, we apply the variational QSVD algorithm by introducing an optimization procedure. That is, after preparing the state $|\psi\rangle_{AB}$, we apply two parameterized quantum circuits $U_{A}(\bm{\theta}_{1})$ and $V_{B}(\bm{\theta}_{2})$. The loss function is
@@
where $N_{m}$ is the number of shots, and $p_{x},q_{x}$ are the two binary strings at the $x$-th shot. $\delta(a,b)$ is 0 if $a\neq b$ and 1 if $a=b$. A classical optimizer is employed to optimize the parameters to minimize the loss function.